OpenStudy (loser66):

In the interval [0,1] \(A=\{x|0\leq x\leq1/3\}\) \(B=\{x|1/3\leq x\leq1\}\) \(C=\{x|x=1/3\}\) \(A=\{x|1/2,x,5\}\) Use your intuition to assign values to P(A),P(B),P(C),P(D) Please, help

OpenStudy (anonymous):

gotta use my intuition to read it !

OpenStudy (anonymous):

I'm not sure but I'm hoping this link will be able to help you: https://github.com/juergenmeinecke/EMET1001/blob/master/sets.rst

OpenStudy (anonymous):

i assume there is something more to this question a point is chosen uniformly on \([0,1]\) for example and \[A=\{x|x\leq \frac{1}{3}\}\] then \[P(A)=\frac{1}{3}\] the length of the interval

OpenStudy (loser66):

The net's fault, not mine, hihi

OpenStudy (loser66):

@satellite73 I got you, but how to find P(C)??

OpenStudy (anonymous):

Sorry is this probability or real analysis?

OpenStudy (loser66):

The tittle of book is" Probability and Statistical Inference" by Hogg, Tanis and Zimmerman

OpenStudy (anonymous):

Oh okay. You're given the uniform distribution (which is continuous). The probability of the random variable attaining a point in a continuous distribution is zero.

OpenStudy (loser66):

You mean P(C) =0??

OpenStudy (anonymous):

Yes

OpenStudy (loser66):

And P(D) =1/2, right?

OpenStudy (anonymous):

Correct, the \((1,5)\) part of the interval is zero, so you're left with 1/2.

OpenStudy (perl):

P(A) = 1/3 , P(B) = 2/3 , P(C) = 0 , P(D) = ?

OpenStudy (loser66):

Thank you so much. One more question, I don't see the definition of " probability of the random variable attaining a point ... is zero", on my book. Can you give me link?

OpenStudy (perl):

can you define D again? D = 1/2 <= x <= 5 ?

OpenStudy (perl):

since the universal set here is [0,1] , then it would just be 1/2

OpenStudy (loser66):

D ={ x | 1/2<x<5}

OpenStudy (loser66):

yes, that is my answer too. In the given interval, it is (1/2,1)

OpenStudy (anonymous):

Think of it this way. \[P(c)=P(c\le X\le c)=\int_c^cf(x)~dx=0\]

OpenStudy (loser66):

Got you. :)

OpenStudy (loser66):

Thanks again.

OpenStudy (perl):

you can think of it, each point has the same probability of being picked . and there are an infinite number of points between 0 and 1 . so the probability of picking any specific point in [0,1] is going to 1 / infinity or zero

OpenStudy (anonymous):

You're welcome!

OpenStudy (loser66):

Thanks @perl too. I learn a lot :)

OpenStudy (perl):

what is interesting is that any number you pick has a probability of zero to occur, but you pick a point each time. so saying something has the probability of zero does not mean it is impossible to happen

OpenStudy (perl):

any number you pick had a probability of 0 of occuring before you picked it, but you do pick a number . so probability of zero does not imply it is impossible to happen

OpenStudy (perl):

it just means, given that each point between 0 and 1 is equally likely , it is very very very unlikely you pick a specific point like 0.23457899

OpenStudy (perl):

essentially the probability of picking a decimal that im thinking of between 0 and 1 , im thinking of a number right now, is zero .

OpenStudy (perl):

so ask someone to pick a number between 0 and 1, and then randomly throw a dart at the line [0,1]

OpenStudy (perl):

the probability your dart will land on the number your friend picked is zero (but not impossible)

OpenStudy (loser66):

I understood. :)

OpenStudy (perl):

clearly the dart does land on the line ;)

OpenStudy (perl):

sorry im pedantic. being pedantic is good when you are first learning a new theory

OpenStudy (perl):

it will resolve ambiguities later on

OpenStudy (perl):

questions that may occur

OpenStudy (anonymous):

\[P(C)=0\]

OpenStudy (anonymous):

the point has no length

OpenStudy (loser66):

No, no. I need seeing the problem from different perspective. Thanks for giving the information.

OpenStudy (anonymous):

oh waht @perl said

OpenStudy (loser66):

yup.

OpenStudy (perl):

is it wrong to steal a medal from a guy who has a lot of gold and give it to a not so wealthy person

OpenStudy (perl):

by steal, i mean undo giving a medal and give it to someone else

OpenStudy (perl):

i dont know what this is called, indian giving?